### v3

# Rigid clusters in shear-thickening suspensions: a nonequilibrium critical transition

Abstract
The onset and growth of rigid clusters in a two-dimensional (2D) suspension in shear flow are studied by numerical simulation. The suspension exhibits the lubricated-to-frictional rheology transition, but the key results here are for stresses above the levels that cause extreme shear thickening. At large solid area fraction, $\phi$, below the stress-dependent jamming fraction, we find a critical $\phi_{c}\left(\sigma,\mu\right)$ where $\sigma$ is a dimensionless shear stress and $\mu$ is the interparticle friction coefficient. For $\phi>\phi_{c}$, the proportion of particles in rigid clusters grows sharply, as $f_{\rm rig}\sim\left|\phi-\phi_{c}\right|^{\beta}$ with $\beta=1/8$, as do the fluctuations in the fraction of rigid particles, with a susceptibility measure $\chi_{\rm rig}\sim\left|\phi-\phi_{c}\right|^{-\gamma}$ with $\gamma=7/4$. The system is thus found to exhibit criticality. The results are shown to depend on an effective field $h\left(\mu\right)$, which provides data collapse near $\phi_{c}$ for both $f_{\rm rig}$ and $\chi_{\rm rig}$. This behavior occurs over a range of stresses, with $\phi_{c}\left(\sigma,\mu\right)$ increasing as the stress decreases.
The rheological response of very concentrated, or dense, suspensions results from their proximity to conditions of jamming Boyer et al. (2011); Morris (2020). Discontinuous shear-thickening (DST) in dense suspensions has been explained as being a result of stress-induced frictional contacts at sufficiently large solid volume fraction $\phi$, i.e., sufficiently close to the jamming fraction, $\phi_{\rm J}^{\mu}$ Seto et al. (2013); Wyart and Cates (2014); the superscript indicates the dependence on an interparticle friction coefficient, $\mu$. The transition in flow properties has been related to a cross-over in scaling Wyart and Cates (2014); Ramaswamy et al. (2021); Malbranche et al. (2022) from low stress with lubricated interactions and frictionless jamming at $\phi_{\rm J}^{0}$, to a high-stress contact-dominated regime with jamming at the frictional $\phi_{\rm J}^{\mu}<\phi_{\rm J}^{0}$. The microscale mechanical basis for the transition is explained by the imposed stress overcoming an interparticle repulsive force, of scale $F_{0}$, to rupture the fluid films between particle surfaces, resulting in a change from lubricated to frictional interactions Morris (2018). Thus, the characteristic stress defining the transition is $\sigma_{0}\sim F_{0}/a^{2}$, with $a$ the particle radius. For stresses small relative to $\sigma_{0}$, most particle interactions are lubricated, while for $\tilde{\sigma}\gg\sigma_{0}$ with $\tilde{\sigma}$ the dimensional bulk shear stress, most close-pair interactions are frictional contacts Wyart and Cates (2014); Singh et al. (2018).
In this scenario, dense suspension properties are related to the network of frictional contacts that results in connected structures, or clusters, which make the material more resistive to flow. These stress-induced clusters, and how they resist the flow, are not well-understood. In particular, the statistical mechanical relationship of the local constraints on motion to larger-scale correlations has not been determined. Recent work has, however, analyzed the contact networks found in simulation of shear-thickening suspensions, to determine signatures of the onset of DST at a non-dimensional stress $\sigma=\tilde{\sigma}/\sigma_{0}=O\left(1\right)$ Gameiro et al. (2020); Sedes et al. (2020); Goyal et al. (2024). Here, we focus on the high stress regime, and the manner in which the rigid jammed state is approached, motivated by the recent report of rigid cluster percolation for these conditions van der Naald et al. (2024).
We consider suspensions at stresses $\sigma\gg 1$, well above the shear-thickening transition, focusing on the flow-induced structures that lead to jamming as $\phi$ approaches $\phi_{\rm J}^{\mu}$. Analysis of an order parameter based on the statistics of rigid clusters demonstrates that the suspension becomes unstable to the formation of rigid clusters via a critical transition at $\phi_{c}\left(\sigma,\mu\right)<\phi_{\rm J}^{\mu}$.
Figure 1: (a) Flow curves for 2D suspensions with particle size ratio $a_{\ell}/a_{s}=1.4$, with 50% of the area occupied by each size, and friction coefficient $\mu=100$, at packing fraction $\phi$ ranging from 0.74 to 0.78. (b) Flow-state diagram with critical transition concentration $\phi_{c}\left(\mu,\sigma\right)$. The solid black line passing through the $\phi_{c}\left(\mu,\sigma\right)$ data points represents the critical transition. The solid red line indicates the CST-DST transition and the solid blue line represents the shear jamming fraction, $\phi_{\rm J}^{\mu}\left(\sigma\right)$, with $\phi_{\rm J}^{\mu}\left(\infty\right)\doteq 0.784$. Regions shaded in gray indicate the conditions studied in the present work.
To perform the rigidity analysis rigorously, we study two-dimensional (2D) suspensions, so that here $\phi$ represents the area fraction occupied by particles. Flow under inertialess conditions (Reynolds number of zero) is simulated by applying a well-established method Seto et al. (2013); Mari et al. (2014), which incorporates repulsive and frictional contact forces along with hydrodynamic effects of Stokes drag and short-range lubrication interactions between particles; more details are presented in the Supplemental Material (SM) sup . Shear-thickening phenomenology in 2D agrees with that in three dimensions, as illustrated by flow curves from simulations in Fig. 1a: for the parameters used here, the suspension exhibits continuous shear-thickening (CST) for $\phi<0.76$ and DST at higher $\phi$; to see that DST begins at $\phi=0.76$, note that if the axes are inverted to plot $\sigma(\dot{\gamma})$, this is the lowest $\phi$ at which $\partial\sigma/\partial\dot{\gamma}\rightarrow\infty$. At the largest stress studied, $\sigma=100$, $\phi_{\rm J}^{\mu}\doteq 0.784$ for the friction coefficient of $\mu=100$, as indicated in Fig. 1b. The flow-state diagram, shown in Fig. 1b, is constructed by estimating the frictionless jamming fraction, $\phi_{J}^{0}\doteq 0.843$, and frictional jamming fraction in the high stress limit, $\phi_{\rm J}^{\mu}\doteq 0.784$, for $\mu=100$, by fitting the Wyart-Cates model Wyart and Cates (2014) to the flow curves. The nose in the flow-state diagram, corresponding to the CST-DST transition, is estimated by noting the point of inflection at $\phi\doteq 0.76$ in the flow curves of Fig. 1a. The artificially large $\mu=100$ approaches the limiting situation where all particles in contact with a compressive force between them are constrained to non-sliding motion at the contacts. We consider a range of smaller $\mu$, and also smaller $\sigma$, below.
We use the tools of rigidity percolation formulated by Henkes et al. Henkes et al. (2016) based on the pebble-game (PG) algorithm Jacobs and Thorpe (1996). The PG is named based on association of degrees of freedom, represented by markers called pebbles, with each particle (or each vertex in a graph in the original development Jacobs and Thorpe (1996)). In the present context, in any configuration, each particle initially has three pebbles to represent its three – two translational and one rotational – degrees of freedom. One or two pebbles are moved for each sliding and nonsliding contact, respectively; these are then associated with the contact (or edge in graph language), indicating the respective number of lost degrees of freedom. When there are only three total remaining degrees of freedom for a connected cluster or subgraph, this structure is identified as rigid by the PG. Note that these structures are not truly rigid, as slight particle deformability plays a role in the mechanics. Recent work van der Naald et al. (2024) has shown, using the same simulation method as we apply, that at large stress and sufficiently large $\phi$, clusters identified as rigid by the PG develop and span the system. Here, we consider the same conditions, probing the emergence of rigid clusters as a function of $\phi$. Thus, this work does not consider the DST transition, but addresses the development of these rigid structures at large stress, where contact interactions dominate. This focuses consideration on the roles of friction and solid fraction in the shaded regions of Fig. 1, although we do consider smaller stress to illustrate that the phenomenon does not extend to arbitrarily small $\sigma$.
Figure 2: Rigid clusters. Particles in rigid clusters contributing to $f_{\rm rig}$ (solid red shading), surface particles of rigid clusters (line shaded gray), and nonrigid (unfilled). The simulation shown is at $\phi=0.756$ with $N=2000$ particles, and is periodic in both the horizontal and vertical directions.
Rigidity order parameter – The percolation of rigid clusters has been investigated for jamming of dry grains Ellenbroek et al. (2015); Henkes et al. (2016); Babu et al. (2023), and more recently for flowing suspensions van der Naald et al. (2024). We consider how a sheared suspension becomes unstable to the development of rigid clusters, as identified by the PG. Fig. 2 demonstrates that rigid clusters of varying sizes up to system-spanning are present at $\sigma=100$ and $\phi=0.756$, below $\phi_{\rm J}^{\mu}$ but close to the critical fraction for the onset of rigidity, as we will show. For each sampling, we define the rigid fraction as
$m_{\rm rig}=\frac{1}{N}\sum_{i=1}^{N}n_{i}\ ,$
where $n_{i}=1$ if particle $i$ is in a rigid cluster with all contacts coming from particles within a rigid cluster, and $n_{i}=0$ otherwise; this excludes ‘surface’ particles identified as part of a rigid cluster but having one or more frictional contacts with particles not within a rigid cluster. The behavior described below is qualitatively the same if these surface particles are included; see SM sup .
[…]
$f_{\rm rig}\left(\phi,\sigma\right)=\langle m_{\rm rig}\rangle\equiv\frac{1}{M}\sum_{\alpha=1}^{M}m_{{\rm rig},\alpha}\ .$ (1)
Figure 3: Order parameter $f_{\text{rig}}$ as a function of distance from the critical packing fraction $\phi_{c}$ at $N=2000$ for different values of stress and static friction coefficient $\mu$. The solid black line represents $1.53\,\left[\phi-\phi_{c}\left(\sigma,\mu\right)\right]^{\beta}$ with $\beta=1/8$, as in the Ising 2D model. The inset presents the variation of $\phi_{c}\left(\sigma,\mu\right)$ with static friction coefficient $\mu$. For $\sigma=1$, $f_{\rm rig}\approx 0$ and we have taken $\phi_{c}=\phi_{\text{max}}=0.83$.
Fig. 3 shows that there is a sharp increase in $f_{\rm rig}$ at solid fraction $\phi_{c}\left(\sigma,\mu\right)$, which increases as either of $\sigma$ and $\mu$ decreases, as shown in Fig. 1, and in the inset to Fig. 3. Unless otherwise indicated, $\phi_{c}$ is the value extrapolated to $N\rightarrow\infty$. For $\sigma=1$, a condition below the DST transition where contact interactions are sparse, $f_{\rm rig}\approx 0$. We note that there is a background of small rigid clusters even at significantly lower packing fractions. These clusters at $\phi_{c}-\phi>0.02$ are small, with almost all containing three or fewer particles; the relationship of these small clusters to the friction coefficient is considered below.
It is well-known that critical transitions are distinguished by large fluctuations Cardy (1996). For the sheared frictional suspension, the relevant fluctuations are those of the order parameter, $\delta m_{\text{rig}}\equiv m_{\text{rig}}-f_{\text{rig}}$. We consider the fluctuations of the extensive net rigidity $N\delta m_{\text{rig}}$:
$\chi_{\text{rig}}=N\langle\delta m_{\text{rig}}^{2}\rangle\ .$ (2)
Fig. 4 shows a well-defined peak of $\chi_{\text{rig}}$ at $\phi\approx\phi_{c}$.
Figure 4: Susceptibility $\chi_{\rm rig}$ as a function of distance from the critical packing fraction $\phi_{c}$ at $N=2000$ for different stresses and static friction coefficient $\mu$.
Figure 5: Scaling collapse of rigid cluster size distribution for a range of packing fraction at stresses $\sigma=$ 10, 20 and 100. The data collapse is shown here for $N=2000$ and static friction coefficient $\mu=100$. The inset shows the average cluster size $\langle n\rangle$ as a function of $\left(\phi-\phi_{c}\right)$ for different stress values.
[…]
Fig. 3 demonstrates that the transition becomes sharper, implying that the noted small clusters for $\phi<\phi_{c}$ occupy a smaller fraction of the system, as $\mu$ decreases. This leads us to the hypothesis that the friction coefficient controls an effective field, $h$, that is related to the background of small clusters. Established scaling theory Cardy (1996) can be used to deduce a form for the order parameter $f_{\rm rig}$ and its fluctuations $\chi_{\rm rig}$, at a given system size, $N\sim L^{2}$, that accounts for the effects of a field:
$\begin{split}\left(\left|h\right|\right)^{-\beta/\Delta}f_{\text{rig}}\left(\delta\phi,h\right)&\propto\mathcal{F}_{N}\left(\frac{h^{1/\Delta}}{\delta\phi}\right)\ ,\\ \left(\left|h\right|\right)^{\gamma/\Delta}\chi_{\text{rig}}\left(\delta\phi,h\right)&\propto\mathcal{F}_{\chi,N}\left(\frac{h^{1/\Delta}}{\delta\phi}\right)\ ,\end{split}$ (3)
where $\delta\phi\equiv\phi_{c}\left(\sigma,\mu\right)-\phi$, and $\beta$, $\nu$, $\gamma$, and $\Delta$ are exponents characterizing the critical point; we emphasize that $\phi_{c}\left(\sigma,\mu\right)$ here denotes the critical packing fraction for the specific system size. There are relationships between these exponents, such as $\Delta=\beta+\gamma$ Cardy (1996).
In Fig. 6, we test our hypothesis by applying the scaling form (Eq. 3) at a fixed $L$ to the data shown in Fig. 3. We find good scaling collapse with exponents $\beta=1/8$, $\gamma=7/4$, and $\Delta=15/8$; these exponents are consistent with those of the 2D Ising model.
Concluding remarks – We have considered dense suspensions that exhibit the lubricated-to-frictional rheological transition Morris (2018), considering only stresses well above the DST transition. Making use of the pebble game algorithm, we identified the rigid clusters as a function of volume fraction at several stresses in this high-stress regime, and we studied their statistics using an order parameter $f_{\rm rig}$, which is the mean fraction of particles in rigid clusters. We find that the suspension becomes unstable to rigid clusters via a critical transition at a volume fraction $\phi_{c}\left(\sigma,\mu\right)$: as $\phi\rightarrow\phi_{c}$ from below, the suspension at large stress becomes unstable to the growth of rigid clusters. This transition is qualitatively illustrated by Fig. 2, showing that there is symmetry between the rigid and nonrigid domains at criticality. We consider slightly deformable particles, and in results not presented, we find that $\phi_{c}$ moves to larger values, i.e., closer to $\phi_{\rm J}^{\mu}$, as the particle hardness increases. The control parameter is the solid fraction $\phi$, and its increase may be seen as analogous to reduction of temperature in a classic model such as the lattice gas, with the jamming fraction $\phi_{\rm J}^{\mu}$ corresponding to $T=0$. The growth of the order parameter as $f_{\rm rig}\sim\left(\phi-\phi_{c}\right)^{\beta}$ with $\beta=1/8$ above the critical solid fraction, and the scaling of the rigidity susceptibility $\chi_{\rm rig}\sim\left|\phi-\phi_{c}\right|^{-\gamma}$, with $\gamma=7/4$, provide excellent descriptions of the data. The evolution of the cluster size distribution, shown in Fig. 5, provides further evidence of criticality. We work well above DST, as indicated in Fig. 1, but have shown that the behavior extends over a range of applied stresses. This implies that there is a line of critical points $\phi_{c}\left(\sigma,\mu\right)$ determined (in a way not yet understood) by the distance from jamming, as the increase of $\phi_{c}$ with decrease of $\sigma$ is consistent with the increase of the jamming fraction $\phi_{\rm J}^{\mu}$ with reduction of stress. Note that $\phi_{c}$ does not extend to arbitrarily small stresses: we find critical behavior at $\sigma=10$, but not at $\sigma=1$, apparently due to the relatively few contacts formed at small stress. Thus, jamming at high stress occurs from a suspension that contains transient rigid clusters, which contrasts with the low stress regime and could provide a natural explanation for the transition in flow properties Wyart and Cates (2014); Ramaswamy et al. (2021); Malbranche et al. (2022). Furthermore, the critical process through which rigid clusters first appear in the flow, promises to provide important insight into the nature of jamming at high stresses.
Acknowledgments
This work was supported by NSF CBET-2228681, CBET-2228680, and DMR-2026834. AS acknowledges Aryabhata HPC facility of IIT (ISM) Dhanbad and National Supercomputing Mission (NSM) for providing computing resources for the HPC System, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. We appreciate helpful conversations with Prof. Emanuela Del Gado, Prof. Silke Henkes, Prof. Heinrich Jaeger, and Dr. Mike van der Naald, to the last of whom we are grateful for assistance in implementing the pebble game algorithm.